3. The basketball just passes over the hands of the attempted blocker and goes into the basket. Determine the shot speed at A, the height of the arc at B, and the ball speed at the basket. Ans: h 11.49 ft 30 YA 7T 10 ft 7 ft

The basketball just passes over the hands of the attempted blocker and goes into the basket. Determine the shot speed at A, the height of the arc at B, and the ball speed at the basket.
Answer #1

motion along horizontal

30 = u cos 30 * t

ut = 34.641... (i)

motion alomg vertical

h = u sin 30 * t - 0.5 * 32.2 t^2

10 - 7 = 0.5* 34.641 - 16.1 t^2

t = 0.943 s

putting kn (i)

u = 36.73 ft/s

=======

b)

time taken by ball to cover 25 ft

t' = 25 / (36.73 * cos 30)

t' = 0.7859 s

again using 2md equation of motion

h - 7 = 36.73* sin 30 * 0.7859 - 0.5 * 32.2*0.7859^2

h = 11.489 ft

=====